+78 nC charge is located at point A (10cm, 10cm ). What is the x component of electric field at point O (0,0)?
Answer #1


electric field at (0,0) is E = k*q/r^2 = (9*10^9*78*10^-9)/(2*0.1^2) = 35100 N/C


X-component of electric field is Ex = -E*cos(45) = -35100*cos(45) = -24819.44 N/C